OpenStudy (anonymous):

Which is the simplified form of 3 times a to the ninth power all over 18 times a cubed.? 1 over 6. a3 1 over 6. a6 1 over 15. a3 15a6

OpenStudy (anonymous):

I feel like it's A

OpenStudy (whpalmer4):

\[\frac{3a^9}{18a^3} = \frac{3}{18}*\frac{a^9}{a^3}\] \[\frac{x^m}{x^n} = x^{m-n}\]

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

I told you everything you need to know to actually find the answer, not just guess...

OpenStudy (anonymous):

I'm not guessing, lol.

OpenStudy (whpalmer4):

\[\frac{3*a*a*a*a*a*a*a*a*a}{18*a*a*a}=\]

OpenStudy (anonymous):

Yeah 9 a's and 4 a's. I'm not sure if the answer is A or B. I know it's one of those for sure though.

OpenStudy (whpalmer4):

I gave you two methods to find the answer. It appears your counting skills might need some work, btw.

OpenStudy (anonymous):

I meant 3. you're pretty rude. Later.

OpenStudy (whpalmer4):

Yep. By the way, my 10 year old just figured out the correct answer.

OpenStudy (anonymous):

That's awesome.